OpenStudy (anonymous):

The price of a pair of headphones is $39.99 before tax. Dalton bought the headphones for 30% off. He then paid 6% sales tax on the discounted price. Part A: How much did Dalton pay in sales tax? Show all work and steps in your solution. Part B: What is the total amount that Dalton paid for the headphones? Show all work and steps in your solution.

OpenStudy (anonymous):

@Vocaloid @ali2x2 @Destinyyyy

OpenStudy (anonymous):

Part A: 39.99 * 0.30 = $12 39.99 - 12.00 = 27.99 Dalton paid $27.99 for taxes.?

OpenStudy (destinyyyy):

Sorry Im trying to solve my own problem.. If you still need help after im done I might be able to help you.

OpenStudy (anonymous):

um ok :)

OpenStudy (destinyyyy):

@Nnesha @whpalmer4

OpenStudy (anonymous):

@QueenofTamaran @sammixboo @campbell_st @ZaneMarcs

OpenStudy (campbell_st):

you calculation is correct for the discounted price... the discount is $12 so he paid $27.99 now find 6% of this amount to get the amount of tax paid.

OpenStudy (campbell_st):

Part B The total amount paid is the discounted price $27.99 + tax... hope it helps

OpenStudy (anonymous):

So i have to find 6% of 27.99?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

would that be 29.70?

OpenStudy (campbell_st):

that's correct.. that will be the tax

OpenStudy (campbell_st):

it should be about $1.80

OpenStudy (anonymous):

i got 1.6794 So round it up to 1.70?

OpenStudy (anonymous):

then that would be 29.69 right?

OpenStudy (anonymous):

then that would be 29.69 right?

OpenStudy (campbell_st):

ok... that makes sense so the total paid is $27.99 + 1.68 = I think your rounding of 1.6794 sohould be $1.68

OpenStudy (anonymous):

I turned in my test already lol

OpenStudy (anonymous):

But thank you :)